OpenStudy (anonymous):

be sure to distribute and collect like terms before you solve 2(4x+3)-9=45 x=___________

OpenStudy (anonymous):

8x + 6 - 9 = 45 8x - 3 = 45 8x = 48 48 divided by 8 = 6. x = 6

OpenStudy (anonymous):

6 - 9 = -3 45 + 3 = 48

OpenStudy (anonymous):

hey you where u from ?

OpenStudy (anonymous):

Me?

OpenStudy (anonymous):

yeah you ?

OpenStudy (anonymous):

IL

OpenStudy (anonymous):

where ?

OpenStudy (anonymous):

Illinois. USA

OpenStudy (anonymous):

ah that's cool, may i have your facebook name ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

why not ?